OpenStudy (rayman5766):

The price of a mobile phone decreased from $375 to $250 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 32% B. 33% C. 50% D. 83%

OpenStudy (anonymous):

Do you know how to find a percent?

OpenStudy (rayman5766):

yes

OpenStudy (rayman5766):

@Data_LG2 help me pls pls pls pls please

OpenStudy (anonymous):

okay so subtract the smaller num from the bigger num, and the divide the result by 100

OpenStudy (anonymous):

^do what she says :)

OpenStudy (rayman5766):

I am let me put in my answer - 1.25

OpenStudy (rayman5766):

were she go did I get it wrong that she just gave up on my like my mom did (crying)

OpenStudy (rayman5766):

no I am just kidding

OpenStudy (rayman5766):

but were did she go

OpenStudy (rayman5766):

@iGreen help me every one around has frozen not the movie but I need help though

OpenStudy (rayman5766):

@Squirrels help me pls pls pls please

OpenStudy (kendricklamar2014):

A. 375 - 32% = $255 B. 375 - 33% = $251 C. 375 - 50% = $188 D. 375 - 83% = $64 What do you think the answer is @rayman5766

OpenStudy (rayman5766):

B

OpenStudy (kendricklamar2014):

Correct

OpenStudy (rayman5766):

thamks man with out you I would been stuck on that all day pluse my mom would get mad and say hurry up like she is doing now great

OpenStudy (kendricklamar2014):

No Problem